# TopLoc: A Locality Sensitive Hashing Scheme for Trustless Verifiable Inference

### v1

@@ 6.1 FP8 Inference and KV-cache compression
Although, our preliminary experiments show that it is possible to distinguish between batches of generations that were done using fp8 vs bf16, the margin between them is small. It thus might only be possible to reliably distinguish them when the device configuration and attention implementation is the same. It also might be necessary to predict how unstable a generation will be based on the validators computation to determine a generation specific dynamic threshold for acceptance.
In this work, we also do not test whether our method is able to distinguish the types of KV-cache compression being used by the inference provider.
We leave these experiments and threshold tuning to future works.
@@ 6.2 Speculative decoding and sampling
Our method is not capable of detecting speculative decoding, a scenario where a provider uses a cheaper model for decoding while relying on the larger model only for prefill computations. In such cases, the provider can generate the tokens using the small model and the prefill vectors using the large model, split them into chunks, and calculate hashes to pass the verification process. Addressing this requires inspecting the execution of the sampling algorithm, which lies beyond the scope of this work and is left for future research.
@@ 6.3 Unstable prompt mining
Inference consumers may attempt to exploit the system by mining for prompts that deliberately increase the likelihood of validation failure. For example, one might be able to find an input prompt that causes an increased amount of catastrophic cancellations early in the computation, which can cascade for long generations. Ensuring that the method is resistant to such attacks remains an important consideration for widespread use of TopLoc for verifying inference computation.
@@ 7 Conclusion
In this paper, we addressed the critical need for trust in large language model (LLM) inference by introducing TopLoc, a novel method for verifiable inference. Our approach tackles the limitations of existing methods, including cryptographic verification, fingerprinting, and tensor activation recording, by significantly reducing storage costs and computational overhead while maintaining robust security guarantees.
TopLoc enables the generation of lightweight, verifiable commits during LLM inference, achieving over 1000x reduction in storage requirements compared to direct tensor recording. It is robust to GPU non-determinism, algebraic reorderings, and diverse inference configurations, ensuring compatibility across varying hardware and execution environments. The method achieves validation speeds up to 100×100$\times$100 × faster than the original inference, making it practical for real-world deployment.
@@
By providing an efficient and reliable foundation for trustless compute protocols, TopLoc advances the usability and transparency of open-source LLM inference. This work opens new opportunities for building decentralized and verifiable AI services, fostering trust in open ecosystems, and enabling broader adoption of open-source models.
@@ B.1 Dataset
A random sample of input prompts from the UltraChat dataset are presented in Table 11.
@@ B.2 Model configurations
Configuration Llama-3.1
8B-Instruct
Llama-3.1
70B-Instruct
INTELLECT-1
Instruct
gemma-2
9b-it
@@ B.3 System prompts
Table 13 shows the system prompts that we used to alter the original user prompt for Section 6.
Table 13: System
Alteration System Prompt

### v2

@@ 6.1 FP8 Inference and KV-cache compression
Although our preliminary experiments show that it is possible to distinguish between generation results that were done using fp8 vs bf16, the margin between them is small. Thus, it might only be possible to reliably distinguish them when the device configuration and attention implementation are the same. It also might be necessary to predict how unstable a generation will be based on the validators computation to determine a generation-specific threshold for acceptance.
In this work, we also do not test whether our method is able to distinguish between the types of KV cache compression being used by the inference provider. We leave these experiments and threshold tuning to future work.
@@ 6.2 Speculative decoding and sampling
Our method is not capable of detecting speculative decoding, a scenario where a provider uses a cheaper model for decoding while relying on the larger model only for prefill computations. In such cases, the provider can generate the tokens using the small model and the prefill vectors using the large model, split them into chunks, and calculate hashes to pass the verification process. Addressing this requires inspecting the execution of the sampling algorithm, which lies beyond the scope of this work.
@@ 6.3 Unstable prompt mining
Inference consumers may attempt to exploit the system by mining for prompts that deliberately increase the likelihood of validation failure. For example, one might be able to find an input prompt that causes an increased amount of catastrophic cancellations early in the computation, which can cascade for long generations. Ensuring that the method is resistant to such attacks remains an important consideration for widespread use of TopLoc and similar methods.
6.4 Spoofing last layer activations
One potential attack vector is to spoof the last hidden layer activations, either by pruning intermediate layers or using a smaller model that mimics the larger model’s activations. However, if a small model is able to reliably reproduce the same layer activations, it effectively means it can match the hidden states of the larger model and thus imply equivalent performance. Given the known capability gap between smaller and larger models, this seems unlikely in practice. However, if the large model is under-trained, this may be a possible attack vector to be explored in future works.
6.5 Difficulty of attack detection
While detecting significant model changes or large prompt alterations using TopLoc is relatively straightforward because of their impact on the top-$k$ activations, more subtle modifications pose greater challenges. In particular, detecting minor prompt tweaks or slight gradient updates in fine-tuned models is significantly harder and requires a higher sensitivity of the detection method. Exploring detection techniques for these subtle changes is an important direction for further research.
@@ 7 Conclusion
In this paper, we address the critical need for trust in large language model (LLM) inference by introducing TopLoc, a novel method for verifiable inference. Our approach tackles the limitations of existing methods, such as cryptographic verification, fingerprinting, and tensor activation recording, by significantly reducing storage costs and computational overhead while maintaining robust security guarantees.
TopLoc enables the generation of lightweight, verifiable proofs during LLM inference, achieving over 1000x reduction in storage requirements compared to direct tensor recording. It is robust to GPU non-determinism, algebraic reorderings, and diverse inference configurations, ensuring compatibility across varying hardware and execution environments. The method achieves validation speeds significantly faster than the original inference, making it practical for real-world deployment.
@@
By providing an efficient and reliable foundation for trustless compute protocols, TopLoc advances the usability and transparency of open LLM inference. This work opens new opportunities for building decentralized and verifiable AI services, fostering trust in open ecosystems, and enabling broader adoption of open models.
Acknowledgements
The authors would like to thank Wei Chun Tan and Benedict Neo for the initial discussion that led to this idea.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Additional Results
Top-k Exact Match Small Deviations Larger Deviations
@@ B.1 Dataset
A random sample of input prompts from the UltraChat dataset is presented in Table 8.
@@ B.2 Model configurations
Configuration Llama-3.1 8B-Instruct Llama-3.1 70B-Instruct INTELLECT-1 Instruct Gemma-2 9b-it
@@ B.3 System prompts
Table 10 shows the system prompts that we used to alter the original user prompt for Section 5.7.
Table 10: System prompts used for prompt alteration experiments
Prompt Alteration System Prompt
@@
Appendix C Subroutines
C.1 Modular polynomial interpolation
Interpolating a polynomial congruence using the Newton’s method has the complexity $O(k^{2})$, where $k$ is the number of top-$k$ values we use in the proof. The pseudocode implementation is provided in Algorithm 3. We also have an open source C++ implementation 555github.com/PrimeIntellect-ai/toploc/blob/main/toploc/C/csrc/ndd.cpp.
The modular inverses required to calculate the congruence can be calculated using the extended Eucledian algorithm, as detailed in Algorithm 4, and has a computational complexity of $O(logM)$, where $M$ is the maximum integer ($2^{16}$ for 16-bit types and $2^{32}$ for 32-bit types).
The value at a point in the polynomial congruence can be calculated using Horner’s method, as described in Algorithm 5 in $O(k)$ time where $k$ is the number of coefficients in the polynomial. In our case, this is the number of top-$k$ values used in the proof.
Algorithm 3 Newton Polynomial Congruence Interpolation
1:  Input: Lists of integers $x,y$ with $|x|=|y|$, Modulus $M$
2:  Output: Coefficients $c$ of interpolated polynomial $P(x)$ in standard form
3:  $n\leftarrow|x|$
4:  $\texttt{dd}\leftarrow y\bmod M$Initializing divided differences}
5:  for $k\leftarrow 1$ to $n-1$ do
6:     for $i\leftarrow n-1$ to $k$ step $-1$ do
7:        $numerator\leftarrow(dd[i]-dd[i-1])\bmod M$
8:        $denominator\leftarrow(x[i]-x[i-k])\bmod M$
9:        $dd[i]\leftarrow(numerator\cdot\texttt{modInverse}(denominator,M))\bmod M$
10:     end for
11:  end for
12:  $c\leftarrow[0]\times n$Output polynomial coefficients}
13:  $\texttt{factor}\leftarrow[1]+[0]\times(n-1)$Rolling factor for polynomial products}
14:  for $i\leftarrow 0$ to $n-1$ do
15:     for $j\leftarrow 0$ to $i$ do
16:        $c[j]\leftarrow(c[j]+dd[i]\cdot\texttt{factor}[j])\bmod M$
17:     end for
18:     if $i+1<n$ then
19:        $m\leftarrow(-x[i])\bmod M$
20:        $prev\leftarrow\texttt{factor}[0]$
21:        $\texttt{factor}[0]\leftarrow(prev\cdot m)\bmod M$
22:        for $k\leftarrow 1$ to $i+1$ do
23:           $temp\leftarrow\texttt{factor}[k]$
24:           $\texttt{factor}[k]\leftarrow(prev+temp\cdot m)\bmod M$
25:           $prev\leftarrow temp$
26:        end for
27:     end if
28:  end for
29:  return $c$
Algorithm 4 Modular Inverse using Extended Euclidean Algorithm
1:  Input: Integers $a$, $m$
2:  Output: Modular inverse of $a$ modulo $m$, if it exists
3:  if $m\leq 1$ then
4:     return 00 {No inverse when modulus is invalid}
5:  end if
6:  $old\_r\leftarrow a$, $r\leftarrow m$Remainders}
7:  $old\_s\leftarrow 1$, $s\leftarrow 0$Bezout coefficients}
8:  while $r\neq 0$ do
9:     $q\leftarrow\lfloor old\_r/r\rfloor$
10:     $tmp\_r\leftarrow old\_r-q\cdot r$
11:     $old\_r\leftarrow r$
12:     $r\leftarrow tmp\_r$
13:     $tmp\_s\leftarrow old\_s-q\cdot s$
14:     $old\_s\leftarrow s$
15:     $s\leftarrow tmp\_s$
16:  end while
17:  if $old\_r\neq 1$ then
18:     throw error {No inverse if $\gcd(a,m)\neq 1$}
19:  end if
20:  return $\texttt{safeMod}(old\_s)$
Algorithm 5 Horner’s Method for Polynomial Evaluation
1:  Input: Coefficients $c=[c_{0},c_{1},\dots,c_{n}]$, Point $x$, Modulus $M$
2:  Output: $P(x)$
3:  $result\leftarrow c_{n}$
4:  for $i\leftarrow 1$ to $n$ do
5:     $result\leftarrow(result\cdot x+c_{n-i})\bmod M$
6:  end for
7:  return $result$
C.2 Injective modulus finding
Finding the injective modulus can be done in $O(k)$ time using the brute force algorithm described in Algorithm 6.
Algorithm 6 Find Injective Modulus
1:  Input: List of integers $x$
2:  Output: Largest modulus $i$ such that $j\bmod i$ is injective for all $j\in x$
3:  for $i\leftarrow 65536$ downto $12^{15}+$ 1 do
4:     $S\leftarrow\{j\bmod i\mid j\in x\}$
5:     if $|S|=|x|$ then
6:        return $i$
7:     end if
8:  end for
9:  raise ValueError("No injective modulus found!")
Although the theoretical worst case constant of $65,536$ can be quite large, on average, the function returns in a few iterations. This is because, assuming the inputs are uniform random, the probability of reaching an iteration decreases exponentially.
Table 11: Distribution of return values from the injective modulus finding function measured with 100 million uniformly random sampled sets of 128 int32 integers
Modulus $m$ Ratio
65536 $0.8833$
65535 $1.031\times 10^{-1}$
65534 $1.203\times 10^{-2}$
65533 $1.401\times 10^{-3}$
65532 $1.607\times 10^{-4}$
65531 $1.927\times 10^{-5}$
65530 $2.192\times 10^{-6}$
65529 $1.818\times 10^{-7}$
65528 $3.030\times 10^{-8}$